Table 13-12
Betty’s Bakery

Quantity
of
cakes


Fixed
Cost


Variable
Cost


Total
Cost

Average
Fixed
Cost

Average
Variable
Cost

Average
Total
Cost


Marginal
Cost

1

$13

$38

2

$28

3

$70

4

$64

5

$110

6

$108

7

$133

8

$185



Refer to Table 13-12. What is the average fixed cost of producing 3 cakes at Betty’s Bakery?

a.

$5.33

b.

$8.33

c.

$2.67

d.

$1.67
Answer #1

The complete table is given below

Q

FC=TC-FC

VC=TC-FC

TC=FC+VC

AFC=FC/Q

AVC=VC/Q

ATC=TC/Q

MC=TC2-TC1

1

25

13

38

25

13

38

2

25

28

53

12.5

14

26.5

15

3

25

45

70

8.333333

15

23.33333

17

4

25

64

89

6.25

16

22.25

19

5

25

85

110

5

17

22

21

6

25

108

133

4.166667

18

22.16667

23

7

25

133

158

3.571429

19

22.57143

25

8

25

160

185

3.125

20

23.125

27

The average fixed cost of producing 3 cakes at Betty’s Bakery is $8.33.

The correct option is : option b.